OpenStudy (anonymous):

Please Help!!! A jet ski depreciates at 11% of its original value each year. If the jet ski was $8000 at its time of purchase, what is the value of the jet ski after 5 years? $13,480.47 $7,120.00 $6,958.86 $4,467.25

OpenStudy (aaronq):

first year the price drops: 8000*0.11=880 New price: 8000-880=7120 2nd year price drop: 7120*0.11=783.2 New price: 7120-783.2=6336.8 can you finish this? or better yet, can you write a function to represent this?

OpenStudy (anonymous):

Nope

OpenStudy (aaronq):

can you finish this with the logic i used?

OpenStudy (anonymous):

Yup, Im not dumb.. I just need a point in the right direction. I don't like the snippy tude.. If u didn't want to help u didn't have to, but the "or better yet, can you write a function to represent this?" was really inappropriate.

OpenStudy (anonymous):

You shouldn't be on a site to help people if u get an attitude doing it. K thanks.. Have a good day. Bye

OpenStudy (aaronq):

i'm not giving you attitude, sorry if that's how i came off. I thought it was more of challenge to ask you to write a function for it.

OpenStudy (anonymous):

Have a nice day.

OpenStudy (condor34):

Actually you are pretty dumb if you think aarong was trying to insult you, obviously you didn't want him to help or point you in the right direction, from what it seems you just wanted the answer. @raebaby420